Determine the distance travelled by the vehicle between time t=120s and time t=160s


1
Expert's answer
2021-09-24T13:25:27-0400



"a=\Delta v/\Delta t=(20-30)/(160-120)=-0.25\ (m/s^2)"


"s=(v^2-v_0^2)/(2a)=(20^2-30^2)/(2\cdot(-0.25))=1000\ (m)" . Answer